OpenStudy (anonymous):

Your friend is trying to calculate the height of a nearby oak tree. You tell him that you learned how to use similar triangles in Geometry class. You tell your friend to measure his height (75 inches) and you measure the length of his shadow (48 inches). Both of you measure the length of the tree’s shadow (38 feet). How tall is the tree (in feet)? Round to the nearest hundredth. Answer

OpenStudy (anonymous):

Get all the units the same so you want feet. The height of your friend is 6.25 feet and the shadow is 4 feet. Then you want do put the equivilant sides on top of each other in a fraction so.\[\frac{ 6.25 }{ x } = \frac{ 4 }{ 38 }\]

OpenStudy (anonymous):

Then multiply it out to get\[4x = 6.25 x 38\] And then divide by 4.

OpenStudy (anonymous):

75 inches equals 6ft 3in. which is 6.25 ft. 48 inches is 4ft 4/6.25 equals 0.64 0.64 multiplied by 38 equals 24.32 24.32 plus 38 equals 62.32 the height of the tree is 62.32 feet.

OpenStudy (anonymous):

\[\frac{ 4 }{ 6.25 } = \frac{ 38 }{x }\]

OpenStudy (anonymous):

That's what you should have done. So multiply both sides by x. Then multiply both sides by 6.25 and divide by 4.

OpenStudy (anonymous):

59.38 is the answer.

OpenStudy (anonymous):

I tried that but it said it was wrong

OpenStudy (anonymous):

Pretty sure that's right.

OpenStudy (anonymous):

Well it said it was wrong

OpenStudy (anonymous):

Maybe you've written the question wrong?

OpenStudy (anonymous):

no I copied and pasted the question